OpenStudy (anonymous):

Which expression has the least value? 3(3 3) 3(3 -3) -3(-3 3) 3(-3 3)

OpenStudy (landshark5):

3(33)= 99 3(3-3)=0 -3(-33)=99 3(-33)= -99 So I would say choices 1 and 3 are tied.

OpenStudy (landshark5):

unless you meant the spaces as multiplication signs.